how do you do 3x+y=-8
-3x+4y=-2?

Answer by checkley77(12844)

3x+y=-8
-3x+4y=-2 add.
----------------
5y=-10
y=-10/5
y=-2 ans.
3x+(-2)=-8
3x-2=-8
3x-8+2
3x=-6
x=-6/3
x=-2 ans.
Proof:
-3*-2+4*-2=-2
6-8=-2
-2=-2